Use Lagrange Multipliers to find the maximum and minimum values of f(x,y)=xy

 subject to the constraint 4x^2+8y^2=16


.


1
Expert's answer
2022-04-25T16:17:40-0400

"f(x,y)=xy, \ \ 4x^2+8y^2=16."

Let "g(x,y)=4x^2+8y^2."


We have "f_x=y,\ \ f_y=x" and "g_x=8x,\ \ g_y=16y."


"\begin{cases}
f_x=\lambda g_x
\\
f_y=\lambda g_y
\\
4x^2+8y^2=16

\end{cases}

\quad 
\begin{cases}
y=8\lambda x
\\
x=16\lambda y
\\
4x^2+8y^2=16
\end{cases}"


"x=16\lambda y=16\lambda \cdot8\lambda x=128\lambda ^2x"

"x\neq 0\ \Rightarrow \ 128\lambda^2=1,\ \ \lambda =\pm \tfrac{1}{8\sqrt{2}}"


"4x^2+8y^2=4x^2+8\cdot (8\lambda x)^2=8x^2=16"

"x=\pm\sqrt{2}," "y=\pm1"


Maximum values: "f(\sqrt{2},1)=f(-\sqrt{2},-1)=\sqrt{2}"

Minimum values:

"f(-\sqrt{2},1)=f(\sqrt{2},-1)=-\sqrt{2}"


Answer: maximum value is "\sqrt{2}" , minimum value is "-\sqrt{2}" .